OpenStudy (anonymous):

a hiker in africa discovers a skull that contains 63% of its original amount of c-14. Find the age of the skull to the nearest year? using the exponential decay formula

OpenStudy (anonymous):

this is so hard to figure out ive tryed like 20 times already :(

OpenStudy (anonymous):

@MDoodler

OpenStudy (anonymous):

Sorry I am at work. You can use the decay formule \( y(t) = a × e^{kt} \) @brittany.hodges Do you know the half-life of C-14 ??

OpenStudy (anonymous):

First find out what the Half-Life of C-14 is, which is 5730. Next, use the decay formula to find out what the rate of decay is. Note, plugin 5730 for t in y(t) and solve for k., which will give you the rate of decay. Once you have the rate, you can use the next equation to get how old the skull is. Decay formula = \( y(t) = a * e^{kt}\) Now use this equation \( t = \frac{ln(\frac{N}{N_0})}{k} * HalfLife\) to find how old the skull is. To use this formula, lets say the half life was 4555 and our skull contained 63% of it c-14 and the rate of decay was -0.693, we would set up the equation like so \(t = \frac{ln(\frac{.63}{100})}{-0.693} * 4555\) If you need more help, type @ and then my name MDoodler and post it here @MDoodler